OpenStudy (anonymous):

The formula for the perimeter of a rectangle is P = 2l + 2w. If P = 104 and l = 13, what is the length of w?

OpenStudy (phi):

This problem requires you to solve an equation. In P= 2L + 2W replace P with 104, and replace L with 13. can you do this? Can you solve the equation?

OpenStudy (anonymous):

w=39?

OpenStudy (phi):

Let's check: P = 2*L + 2*W, with L= 13 and W=39 P= 2*13 + 2*39 P= 26+ 78 P= 104 which matches the question. Looks good!